OpenStudy (anonymous):

two tangents are drawn to circle O from external point P. If the major arc formed has a measure of 280, find m

jimthompson5910 (jim_thompson5910):

Which one do you think it is?

OpenStudy (anonymous):

140 but im not sure at all

jimthompson5910 (jim_thompson5910):

If the major arc is 280, then what is the minor arc?

OpenStudy (anonymous):

Is it half the major arc?

jimthompson5910 (jim_thompson5910):

no, the two arcs add to 360 degrees (since they make up a full circle)

jimthompson5910 (jim_thompson5910):

so major arc + minor arc = 360 and minor arc = 360 - major arc

OpenStudy (anonymous):

so the minor arc is 80 right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

which would make m<p 80 since its a central angle?

jimthompson5910 (jim_thompson5910):

so the central angle that subtends the tangent points is ?

jimthompson5910 (jim_thompson5910):

no not quite

jimthompson5910 (jim_thompson5910):

the central angle is 80, but m < P is not 80

OpenStudy (anonymous):

hm im not sure...it cant be half of 80

jimthompson5910 (jim_thompson5910):

if you draw a picture, you'll see that a quadrilateral forms Two angles of this quadrilateral are 90 degrees (since they are the tangent points). So we have 90+90 = 180 degrees so far. The remaining two angles MUST add to 180 degrees because all four angles in any quadrilateral add to 360 degrees. So central angle + < P = 180

jimthompson5910 (jim_thompson5910):

So < P = 180 - 80 = 100

OpenStudy (anonymous):

so <p is 100?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

thank you so much for actually explaining it to me!

jimthompson5910 (jim_thompson5910):

you're welcome